OpenStudy (anonymous):

tan^2 x = 1-cos2x/ 1+ cos2x true or false?

OpenStudy (anonymous):

That would be true. It' an identity.

OpenStudy (anonymous):

It's not an identity. Does tan^2x = 1-(cos2x/1)? I'm certain it doesn't. It does, however, equal to (1-cos2x)/2. This is an example of a half-angle identity.

OpenStudy (anonymous):

er....er

OpenStudy (anonymous):

\[\frac{ 1-\cos 2x }{ 1+\cos 2x }=\frac{ 1-(1-2\sin^2x) }{ 1+(2\cos^2x-1) }=\frac{ 2\sin^2x }{2\cos^2x }=\tan^2x\]

OpenStudy (anonymous):

so its true.

OpenStudy (anonymous):

um 0.0 :/

OpenStudy (anonymous):

something wrong?

OpenStudy (anonymous):

you guys are confusting me 0.o true or no?

OpenStudy (anonymous):

the first i did was to use the double angle formula for sine and cosine. then it was simplified, then since the ratio of sine and cosine is tangent, then the next step is obvious.